OpenStudy (anonymous):

What is the simplified form of the expression? x^8-2y^10-5x^5

OpenStudy (anonymous):

I see nothing too simplify, It stays the same ! IMAO :(

OpenStudy (anonymous):

that ^ sign is for pwers

OpenStudy (anonymous):

\[\Large x^8-2y^{10}-5x^5\] if this is what you mean by ^ ... then my answer is still the same !

OpenStudy (anonymous):

A:10x^80y^10 B:10xy^23 C:8x^13y^10 D:10x^13y^10 choices

OpenStudy (anonymous):

there's now way of simplifying this EXCEPT if you made a mistake ... and wrote MINUS instead of MULTIPLY ... with minus and plus this stays the same... if you have to multiply, then REWRITE your expression... DRAW it ... ;) otherwise, I can't help you ...

OpenStudy (anonymous):

there's no way of ....

OpenStudy (anonymous):

yea its multiply sorry

OpenStudy (anonymous):

is it ..\[\Large x^8\cdot 2y^{10}\cdot 5x^5\] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so.. then :\[\LARGE 2\cdot 5 \cdot x^{8+5}\cdot y^{10}=10x^{13}\cdot y^{10}\]

OpenStudy (anonymous):

Remember these... \[\LARGE a^m\cdot a^n=a^{m+n}\] \[\LARGE \frac{a^m}{a^n}=a^{m-n}\] \[\LARGE (a^m)^n=a^{m\cdot n}\] \[\LARGE \sqrt[m]{a^n}=a^{\frac{n}{m}}\] @RYSSA2K10

OpenStudy (anonymous):

did you get the answer?

OpenStudy (anonymous):

(x^8)(2y^10)(5x^5)

OpenStudy (anonymous):

rearranging them 2(5)(x^5)(x^8)(y^10) =10x^13 y^10.....answer D

OpenStudy (anonymous):

what is the simplified form of the expression [m^-1 m^5/m^-2]^3